Hayhayz (hayhayz):

Remember to identify all your data, write the equation, and show your work. King Kong carries Fay Wray up the 321-meter high Empire State Building. At the top of the skyscraper, Fay Wray’s shoe falls from her foot. Neglecting air resistance, how long does it take for the shoe to reach the ground?

Hayhayz (hayhayz):

@AloneS

Hayhayz (hayhayz):

@imqwerty

imqwerty (imqwerty):

Use this equation- S=ut+ (1/2)at^2 S-displacement of particle u-initial velocity of particle a-acceleration of particle t-time taken

imqwerty (imqwerty):

\[S=ut+\frac{1}{2}at^2\]

imqwerty (imqwerty):

The shoe falls off from her foot so this means that the shoe didn't had any initial velocity or we can say that the initial velocity was 0 because the shoe just dropped off So u=0

imqwerty (imqwerty):

Now here gravity is accelerating the shoe And gravitational acceleration is equal to 9.8m/s^2 so a=9.8

imqwerty (imqwerty):

The shoe is falling from the top of the building to the ground so its displacement is equal to the height of the building So S=321

imqwerty (imqwerty):

Now we jus gotta plug in the stuff and get t S=ut+(1/2)at^2 321=0*t + (1/2)*9.8*t^2 321=4.9t^2 From here u can get the time "t"

Hayhayz (hayhayz):

Thank you so much qwerty! <3

imqwerty (imqwerty):

Np<3(: